What charge is stored in a 200uF capacitor when 100 v is applied to it?
1
Expert's answer
2018-02-05T08:25:07-0500
We know that
Q= CV
= 200 ×10^-6 × 100 C
= 20000 ×10 ^-6 C
= 2×10^-2 Coulomb